QUESTION 7

  1. In developing patient appointment schedules, a medical center wants to estimate the mean time that a staff member spends with each patient. How large a sample should be taken if the desired margin of error is two minutes at a 99% level of confidence? Use a planning value for the population standard deviation of 8 minutes.

    100

    102

    104

    106
Answer #1

Population standard deviation, \small \sigma = 8

Margin of error, E = 2

At \small \alpha = 1-0.99 = 0.01, critical value, a/2 = ABS(NORM.S.INV(0.01/2)) = 2.576

Sample size =

rl 30 2.576 8 2 106.17 106